# Not all Layers of LLMs are Necessary during Inference

### v2

@@ Acknowledgments
We utilized the sklearn library for training SVM111https://scikit-learn.org/stable/modules/svm.html and CRF222https://sklearn-crfsuite.readthedocs.io/en/latest/, adhering to default configurations. For SVM and CRF training, we used the sklearn library with default settings. Given a training dataset with training examples, the time complexity for SVM training typically ranges from to , where is the feature dimension. SVM prediction time complexity is per single inference. For standard linear-chain CRF, the training time complexity is approximately , where is the average sequence length, is the label count. The prediction time complexity for CRF is per single inference. In contrast, the inference time complexity for large models like llama2 is per single inference, where is the hidden size, is the sequence length, and represents the number of layers. Comparatively, the computational load of SVM and CRF is negligible when compared to large models.
@@
Given a language model with transformer layers, hidden size , sequence length , vocabulary size , and batch size . Each transformer block needs FLOPs for the forward pass. The other main contributor to the FLOPs count is the classification layer in the language model head, which transforms features of dimension to the vocabulary dimension . The required FLOPs for this operation is in the forward pass. While AdaInfer does convert hidden states to logits at each block through classification layer, it only utilizes the hidden state from the last token for conversion, even when the sequence length is 2048 or longer. In the case of Llama2 7/13/70B, this computation accounts for only 0.000288, 0.000236, and 0.000152 of the total number of FLOPs for transformer inference. Similarly, for OPT 13B, it amounts to 0.000367. Consequently, the computational burden associated with this aspect can be disregarded. Summing these together, a transformer model with transformer layers, the total number of floating-point operations for inference is . Thus, the ratio of inference cost in FLOPs can be calculated through Equation 2.
Appendix B Details of Creating Training Data for Classifier
Considering a training input instance and its corresponding label from . Once is processed through a decoder layer of LLM, we can extract a general feature vector ( is the number of features). Additionally, we obtain the probability distribution over the vocabulary of the current layer’s hidden state after passing through the classification layer (as depicted in Section 3.1). This can be represented as: , where is the hidden state of the current layer, and are the weights and bias of the classification layer, respectively. Function is applied to convert logits to probabilities. Let the highest-ranked token in this distribution be denoted as , where finds the token with the highest probability. If matches the label , the associated label for the feature vector is designated as positive; otherwise, it is labeled as negative. Thus, for an layer LLM, each input instance yields pairs of .
@@ Appendix D Comprehensive Summary of Results
The results of all LLMs using different classifiers are summarized in Table 8 and 9. We have highlighted the best results for each task in the current setting. The experimental results indicate that (i) early exits are feasible for different tasks, (ii) the timing of early exits varies depending on the instance, and (iii) in both zero-shot and few-shot settings, accuracy is comparable with baseline models. It’s worth noting that for individual tasks, AdaInfer even outperforms the baseline in zero-shot or few-shot accuracy. This suggests that in inference scenarios, deep layers may tend to over-represent some tasks, potentially impairing performance.

### v3

@@ Acknowledgments
We utilized the sklearn library for training SVM333https://scikit-learn.org/stable/modules/svm.html and CRF444https://sklearn-crfsuite.readthedocs.io/en/latest/, adhering to their default configurations. Given a training dataset with training examples, the time complexity for SVM training typically ranges from to , where is the feature dimension. SVM prediction time complexity is per single inference. For standard linear-chain CRF, the training time complexity is approximately , where is the average sequence length, is the label count. The prediction time complexity for CRF is per single inference. In contrast, the inference time complexity for large models like llama2 is per single inference, where is the hidden size, is the sequence length, and represents the number of layers. Comparatively, the computational load of SVM and CRF is negligible when compared to large models.
@@
Given a language model with transformer layers, hidden size , sequence length , vocabulary size , and batch size , each transformer block needs FLOPs for a forward pass. The other main contributor to the FLOPs count is the classification layer in the language model head, which transforms features of dimension to the vocabulary dimension . The required FLOPs for this operation is in the forward pass. While AdaInfer does convert hidden states to logits at each block through classification layer, it only utilizes the hidden state from the last token for conversion, even when the sequence length is 2048 or longer. In the case of Llama2 7/13/70B, this computation accounts for only 0.000288, 0.000236, and 0.000152 of the total number of FLOPs for transformer inference. Similarly, for OPT 13B, it amounts to 0.000367. Consequently, the computational burden associated with this aspect can be disregarded. Summing these together, a transformer model with transformer layers, the total number of floating-point operations for inference is . Thus, the ratio of inference cost in FLOPs can be calculated as
Appendix B Classifier Training
Considering a training input instance and its corresponding label from . Once is processed through a decoder layer of LLM, we can extract a general feature vector ( is the number of features). Additionally, we obtain the probability distribution over the vocabulary of the current layer’s hidden state after passing through the classification layer (as depicted in Section 3.1). This can be represented as: , where is the hidden state of the current layer, and are the weights and bias of the classification layer, respectively. Function is applied to convert logits to probabilities. Let the highest-ranked token in this distribution be , where finds the token with the highest probability. If matches the label , the associated label for the feature vector is designated as positive; otherwise, it is labeled as negative. Thus, for an layer LLM, each input instance yields pairs of .
@@ Appendix D Comprehensive Summary of Results
The results of all LLMs using different classifiers are summarized in Tables 7 and 8. We highlight the best results for each task in the current setting.555We report the top-1 accuracy score on the test set following function vectors Todd et al. (2024) (HELM implementation). The experimental results indicate that (i) early exits are feasible for different tasks, (ii) the timing of early exits varies depending on the instance, and (iii) in both zero-shot and few-shot settings, accuracy is comparable with dense models. It’s worth noting that for individual tasks, AdaInfer even outperforms the dense model in zero-shot or few-shot accuracy. This suggests that in inference scenarios, deep layers may tend to over-represent some tasks, potentially impairing performance.